OpenStudy (anonymous):

how to solve 5x-12=2x+6

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

5x-12=2x+6 +12 +12 5x=2x+18 -2 -2 3x=18 divide by 3 and you will have x